OpenStudy (bananas1234):

For the circle with equation (x – 2)^2 + (y – 1)^2 = 16 2. Answer each question. (a) What are the coordinates of the center? (b) What are the radius and diameter of the circle? (c) Graph the circle.

OpenStudy (bananas1234):

@Directrix @Nnesha

OpenStudy (bananas1234):

@zepdrix

OpenStudy (mathmale):

It's important that you know and understand how to use the general equation of a circle of radius r, centered at the point (h,k):\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (mathmale):

Compare your particular equation to this one. What are y our h, k and r?

OpenStudy (bananas1234):

k is 1 h is 2 and y is unknown? @mathmale

OpenStudy (mathmale):

Stop for a minute and think of what you're trying to do: Find the center and radius of a given equation of a circle. You were not asked to find x or y; you were asked to find the coordinates of the center of the circle, represented by (h,k). So, it doesn't help if you write "y is unknown." What is h? k? r? You were given: (x – 2)^2 + (y – 1)^2 = 16 Please write the general equation of a circle centered at (h,k) with radius 2 below the given equation. Compare them. Find h, k and r thru this comparison.

OpenStudy (bananas1234):

(x – 2)^2 + (y – 1)^2 = 16 (x – h)2 + (y – k)2 = r2 H = 2 K = 1 idk how to find r @mathmale

OpenStudy (bananas1234):

@mathmale are you still there?

OpenStudy (bananas1234):

@zepdrix can you help me finish this? @mathmale went offline

zepdrix (zepdrix):

Sup? Where we at? :o

OpenStudy (bananas1234):

read above

OpenStudy (bananas1234):

Im so lost this was the most confusing unit of math ever

zepdrix (zepdrix):

\[\large\rm (x-2)^2+(y-1)^2=16\]So you've determined that the center is located at \(\large\rm (h,k)=(2,1)\). And next we need to figure out the radius.

OpenStudy (bananas1234):

ok

zepdrix (zepdrix):

The right side of the equation is supposed to be written as a square. It's not very helpful when it's not written in that form. Is there some way to rewrite 16 as a number squared?

OpenStudy (bananas1234):

4^2

zepdrix (zepdrix):

\[\large\rm (x-2)^2+(y-1)^2=4^2\]Ooo that might help us, that more closely resembles the general form of a circle, ya?\[\large\rm (x-h)^2+(y-k)^2=r^2\]

zepdrix (zepdrix):

Are you able to spot the radius when it's in that form?

OpenStudy (bananas1234):

4

zepdrix (zepdrix):

Good good good. Understand how to get the diameter from that information?

OpenStudy (bananas1234):

im not quite sure 4 times 2 = 8, is that right?

zepdrix (zepdrix):

yay good job!

OpenStudy (bananas1234):

ok whats next?

zepdrix (zepdrix):

we'll use these key features that we've determined so far, to help us graph the circle.

OpenStudy (bananas1234):

ok

zepdrix (zepdrix):

In the upper right corner of the picture, there is a "copy picture" button. Do you know how to use that? Wanna try labeling the center?

OpenStudy (bananas1234):

sure i think i can do that

OpenStudy (bananas1234):

 is that correct?